OpenStudy (anonymous):

cos^2(pi/8+x/2) - sin^2(pi/8-x/2) is equal to

ganeshie8 (ganeshie8):

use this : \(\large \cos(\theta) = \sin(\frac{\pi}{2} - \theta)\)

OpenStudy (anonymous):

sin(pi/2-pi/8-x/2)

ganeshie8 (ganeshie8):

yes :) \[\large \cos\left(\frac{\pi}{8} + \frac{x}{2}\right) =\sin\left(\frac{\pi}{2} - \left(\frac{\pi}{8} + \frac{x}{2}\right)\right) \]

ganeshie8 (ganeshie8):

simplify pi/2-pi/8

OpenStudy (anonymous):

3pi/8

ganeshie8 (ganeshie8):

Ahh yeah it doesn't help :/

ganeshie8 (ganeshie8):

try double angle identities : \[\large \cos^2 \theta = \dfrac{1+\cos2\theta }{2}\] \[\large \sin^2 \theta = \dfrac{1-\cos2\theta }{2}\]

ganeshie8 (ganeshie8):

\[ \large \cos^2(\pi/8+x/2) - \sin^2(\pi/8-x/2)\] \[ \large \dfrac{1+\cos(\pi/4+x)}{2} - \dfrac{1-\cos(\pi/4-x)}{2} \] \[ \large \dfrac{1}{2}(~ \cos(\pi/4+x) + \cos(\pi/4-x) ~) \] \[ \large \dfrac{1}{2}(~ 2\cos(\pi/4) \cos(x)~) \] \[ \large\cos(\pi/4) \cos(x) \]

OpenStudy (anonymous):

1/sqrt(2)cosx

OpenStudy (anonymous):

yes @No.name

OpenStudy (anonymous):

Is theta the same or different in both the cases would you check it please

OpenStudy (anonymous):

different

OpenStudy (anonymous):

oh o k